OpenStudy (anonymous):

The law of cosines can only be applied to nonright triangles. True of false.

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Actually, the law of cosines for right triangles is a pretty cool thing to check out.

OpenStudy (anonymous):

The law of cosines is: c^2 = a^2 + b^2 - 2abCos(C) Where C is the angle opposite from the side c. So let's see what happens when C is 90 degrees. 

OpenStudy (anonymous):

We get c^2 = a^2 + b^2 -2abCos(90) What is cos(90)? Well, it's 0. So this becomes: c^2 = a^2 + b^2 - 2*ab*0 2*ab*0 = 0, so c^2 = a^2 + b^2 which is the Pythagorean theorem! So the Pythagorean theorem which you have known for so long is actually a specific case of the law of cosines! Super cool! =D